the acceleration due to gravity on the planet Calcllceara is -10m/s^2. A stone is thrown downwards with a speed of 5 m/s from the upper observation deck of the Base Camp Alpha Tower, 150 m above the ground
1
Expert's answer
2020-12-22T18:35:42-0500

Kinematic equations for objects thrown downwards

We assume here that acceleration equals "-a=-10 m\/s^2"  (with the positive direction upward).


"v=v_0+at"

"y=y_0+v_0t+\dfrac{at^2}{2}"


"v^2=v_0^2+2a(y-y_0)"



Given "y_0=150m, v_0=-5m\/s, a=-10m\/s^2"



"y_1=0:0=150-5t-\dfrac{10t^2}{2}"

"t^2+t-30=0"

"t_1=-6, t_2=5"

Since "t\geq0," we take "t_2=5s."


"v(5)=-5-10(5)=-55(m\/s)"

The stone will fall down the ground after 5 seconds. Its velocity will be 55 m/s when it falls down the ground.